# Map2Text: New Content Generation from Low-Dimensional Visualizations

### v1

@@ 3.3. Definition
Moderate The atomic statement can be logically inferred from the reference, or the atomic statement can support the reference.
@@
Table 1. Overview of different strictness criteria. A bolded letter of each level is used as its abbreviation in later references. Details are provided in Appendix B.
Comment
Atometric builds on the strengths of existing evaluation metrics while addressing their limitations. Atometric Precision and Atometric Recall can be viewed as relaxations of BLEU and ROUGE, respectively, as they shift from strict n-gram matching to statement-level alignment. This allows generation with more flexible expression while preserving underlining semantic and logical relationships. Compared to BERTScore and BLEURT, which rely on embedding similarity, Atometric uses LLM-based multi-level relevance judgments, offering a more interpretable and adjustable evaluation. By focusing on atomic statements and incorporating varying strictness levels, Atometric provides a more nuanced and detailed assessment of content generation, particularly suited to tasks like Map2Text that require a balance between semantic accuracy and logical coherence.
@@ 4. Experiments
With the new evaluation metric, we are able to benchmark a set of simple yet effective candidate methods for Map2Text, instantiating the outlined solutions in Section 2.2, and compare them with a human baseline and a strong dummy baseline.
@@ 4.1. Candidate Methods
Fine-Tuning (FT)
The most straightforward approach is to directly learn the inverse mapping $g:{\mathbb{R}}^{2}\to{\mathcal{S}}$ by fine-tuning a pre-trained model. We fine-tune the 70B Llama 3.1 model using LoRA (Dubey et al., 2024), where the model takes a 2D coordinate as input and outputs the corresponding text. For instance, an input query would be, ”Convert the coordinate to text: [9.1054, 10.1339].”
Embedding Inversion
If high-dimensional embeddings are available, we can decompose $g$ into two separate mappings: $g_{1}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{H}$ and $g_{2}:{\mathbb{R}}^{H}\to{\mathcal{S}}$. For the first mapping, $g_{1}$, we use a k-NN interpolation algorithm to project the 2D coordinates into the $H$-dimensional embedding space. While more complex methods could be employed for $g_{1}$ (e.g., training a neural network), we select this straightforward approach as it requires no additional training and demonstrates comparable empirical performance. The second mapping, $g_{2}$, then converts the high-dimensional embeddings into text using the vec2text model (Morris et al., 2023). Since we use the pre-trained vec2text model without any additional fine-tuning, it is limited to supporting embeddings generated specifically by text-embedding-ada-002.
Retrieval-Augmented Generation (RAG)
Prompting LLMs has shown strong performance across various tasks (Lewis et al., 2020). However, a straightforward zero-shot prompting approach—like the query mentioned in the fine-tuning example above—is not feasible as a pre-trained LLM lacks inherent knowledge of the visualization map. To address this limitation, we incorporate a set of Retrieval-Augmented Generation (RAG) methods on GPT-4o to provide relevant context and enhance the model’s understanding of the map. We use the same notation in Section 2 to illustrate each method. For each query position ${\bm{x}}_{q}\in{\mathbb{R}}^{2}$,
@@
    (1st-order RAG) We retrieve the texts corresponding to its nearest neighbors $\mathcal{N}(\mathbf{x}_{q})\subset\{1,\dots,N\}$ and use this information to prompt the LLM to generate new content $\hat{s}$.
@@
    (1st-order RAG with few-shot examples) Building on (1), we add few-shot examples. Each few-shot example consists of texts from $\mathcal{N}(\mathbf{x}_{i})$ and the corresponding $s_{i}$ for some $i\in\mathcal{I}$, where $\mathcal{I}$ is a pre-selected set.
@@
    (2nd-order RAG with few-shot examples) Building on (2), we replace the pre-selected set $\mathcal{I}$ with $\mathcal{N}(\mathbf{x}_{q})$. Intuitively, this means we consider information from ”2-hop” neighbors of the query position.
@@
    (1st-order RAG with chain-of-thought examples) Building on (2), we add additional chain-of-thought reasoning (Wei et al., 2022) as auxiliary information.
Implementation of each method can be found in Appendix D.
4.2. Datasets
We evaluate the candidate models on four diverse datasets to demonstrate the broad application of Map2Text.
Dataset Name # Entries Avg. Len. Vis. Pipeline
Persona 100k 27.10 gte-v1.5 + UMAP
Red-Teaming Strategies 39k 84.06 ada-002 + UMAP
CS Research Idea 149k 41.57 ada-002 + LargeVis
CS Research Context 150k 40.96 ada-002 + LargeVis
Table 2. Basic information of datasets and visualizations. Average length is calculated using tiktoken. A bolded letter of each dataset is used as its abbreviation in later references.
Persona
We use an existing visualization [^3] https://huggingface.co/datasets/argilla/FinePersonas-v0.1-clustering-100k of a subset of the PersonHub dataset (Chan et al., 2024) where each text entry provides a brief description of a synthetic persona. Map2Text can be applied to generate new persona descriptions based on different regions of the map, enabling the creation of diverse character profiles.
Red-Teaming Strategies
We construct a summarization dataset [^4] Details of the construction are provided in Appendix C.2. from the LLM Red Teaming dataset by Ganguli et al. (2022) which comprises human-generated and annotated red-teaming dialogues aimed at identifying vulnerabilities in LLMs. Given a 2D visulization of these summaries, Map2Text can help generate unexplored red-teaming strategies, enhancing the robustness testing of LLMs.
CS Research Ideas & Context
We create 2D maps of the MASSW dataset (Zhang et al., 2024), where each entry corresponds to the research context or the key idea of a computer science publication. A good Map2Text model can assist scientific innovation by generating new research ideas or new context of emerging topics from unexplored regions on the map, helping researchers to identify gaps in current research and inspire novel directions for scientific inquiry.
The statistics of these datasets are summarized into Table 2. Note that the visualization pipelines for these datasets are different and are hidden from candidate Map2Text methods (except for embedding inversion), and a robust model should be capable of generalizing across different configurations.
@@ 4.3. Experiment Setup
Each dataset is divided into training and testing sets, as described in Section 2.3, with 200 samples consistently held out for testing across all datasets. We apply a random split for the Persona and Red-Teaming Strategies datasets and a time-based split for the CS Research Ideas & Context datasets, where the most recent 200 entries (all in Dec 2023) are held out for testing, simulating the scenario of real-world research exploration.
@@
We prioritize our proposed evaluation metric, Atometric , as it is specifically designed to capture the nuanced semantic and contextual alignment. We report Atometric-F1 across three levels of strictness: loose (L), moderate (M), and strict (S), as well as Atometric-Precision (M) and Atometric-Recall (R) for a more comprehensive picture [^5] We use gpt-4o-2024-05-13 as the backbone LLM in Atometric.. In addition to Atometric, we include traditional evaluation metrics—BERTScore (F1) (Zhang et al., 2019), BLEURT (Sellam et al., 2020), METEOR (Lavie and Agarwal, 2007), and ROUGE-2 (Lin, 2004)—as reference points for comparison. Evaluation on some extra metrics such as BLEU and other Atometric scores are provided in Appendix A.1.
@@ 4.4. Baselines
We recruit human annotators to conduct the Map2Text task using the Persona dataset, as this dataset does not require special domain expertise. Detailed settings for this study are provided in Appendix A.2. In brief, human annotators are given access to a partial visualization containing the query location and they are allowed to view the textual content of all visible data points. No constraints are imposed regarding their generation strategy, allowing them complete flexibility in navigating the visualization and generating content based on the map. We expect the human baseline to represent a strong benchmark compared to AI methods.
@@
For each query ${\bm{x}}_{q}$, the EchoNearest model simply retrieves and outputs the content of its closest neighbor in the 2D space. This simple baseline operates under the assumption that semantically similar content is clustered closely in the visualization map, making the nearest neighbor a relevant choice for responding to the query, though it lacks novelty. This approach mimics a “lazy” human strategy for content generation which could be effective when novelty is not required. As our gold standard references all exhibit some level of novelty compared to their neighbors, we expect it to perform below both the human baseline and more advanced AI methods when evaluated using proper metrics.
@@
Metric Human EchoNearest
P P R I C
Atometric-F1 (L) 0.901 0.884 0.783 0.430 0.538
Atometric-F1 (M) 0.605 0.556 0.586 0.191 0.105
Atometric-F1 (S) 0.302 0.268 0.282 0.098 0.045
Atometric-P (M) 0.610 0.548 0.598 0.199 0.103
Atometric-R (M) 0.601 0.565 0.575 0.184 0.107
BERTScore F1 0.894 0.898 0.902 0.857 0.862
BLEURT Scores 0.431 0.454 0.436 0.325 0.323
METEOR 0.232 0.306 0.428 0.160 0.170
ROUGE-2 0.110 0.117 0.231 0.032 0.034
The evaluation results for the human and EchoNearest baselines are provided in Table 3. Several observations can be made.
First, under all Atometrics, the human baseline outperforms the EchoNearest model, as expected. However, under conventional metrics, the human baseline scores lower than the EchoNearest model. This discrepancy likely arises because conventional metrics emphasize lexical-level text similarity, favoring the EchoNearest model’s repetition of existing entries. In contrast, human responses, while more contextually relevant, tend to vary in language style and phrasing, which conventional metrics fail to capture. This validates the strength of Atometric in recognizing the semantic and contextual richness of human-generated content.
Second, the four datasets exhibit varying levels of complexity. The EchoNearest model performs relatively well on the Persona and Red-Teaming Strategies datasets, suggesting that the text entries clustered closely in the visualization are highly similar, making these tasks easier to address. In contrast, the model’s performance on the CS Research Idea and CS Research Context datasets is notably lower. This indicates that these tasks require generating more complex or novel content, which demands a deeper understanding of the research domain, beyond simple proximity in the 2D space.
In addition, the Atometric results across different levels of strictness reveal subtle distinctions between research contexts and research ideas, showcasing the strength of Atometric in capturing varying levels of sensitivity. At lower strictness levels, Atometric F1 (L) scores are higher for the CS Research Contexts. While each paper addresses a unique problem, the overarching topics are often related, leading to greater alignment in topic-level relevance. In contrast, at medium and high strictness levels (M and S), research ideas exhibit higher Atometric scores. This occurs because research ideas, while varied, frequently share concrete methodologies or designs, which stricter criteria capture more effectively. By distinguishing these nuanced differences, Atometric demonstrates its ability to assess content sensitivity at multiple levels.
@@ 4.5. Benchmark Results
Persona Atometric-F1 (L) 0.893 0.852 0.824 0.822 0.822 0.783 -
Atometric-F1 (M) 0.602 0.494 0.464 0.454 0.463 0.390 -
Atometric-F1 (S) 0.254 0.183 0.168 0.158 0.194 0.117 -
Atometric-P (M) 0.582 0.460 0.449 0.430 0.459 0.365 -
Atometric-R (M) 0.623 0.533 0.480 0.481 0.467 0.420 -
BERTScore F1 0.898 0.895 0.893 0.892 0.893 0.887 -
BLEURT Scores 0.474 0.454 0.441 0.433 0.406 0.416 -
METEOR 0.329 0.290 0.263 0.260 0.251 0.268 -
ROUGE-2 0.116 0.096 0.078 0.086 0.098 0.074 -
@@
Atometric-F1 (M) 0.514 0.480 0.451 0.414 0.497 0.437 0.428
Atometric-F1 (S) 0.202 0.200 0.168 0.144 0.185 0.167 0.126
Atometric-P (M) 0.479 0.445 0.397 0.358 0.508 0.384 0.329
Atometric-R (M) 0.555 0.521 0.520 0.491 0.486 0.507 0.612
BERTScore F1 0.904 0.902 0.901 0.897 0.894 0.898 0.877
BLEURT Scores 0.446 0.439 0.435 0.417 0.407 0.430 0.394
METEOR 0.433 0.424 0.412 0.371 0.390 0.405 0.312
ROUGE-2 0.250 0.250 0.244 0.232 0.203 0.235 0.114
@@
Atometric-F1 (M) 0.213 0.220 0.214 0.202 0.163 0.210 0.180
Atometric-F1 (S) 0.091 0.098 0.102 0.099 0.084 0.100 0.096
Atometric-P (M) 0.222 0.240 0.272 0.250 0.203 0.263 0.193
Atometric-R (M) 0.205 0.202 0.176 0.169 0.136 0.175 0.169
BERTScore F1 0.860 0.860 0.864 0.864 0.858 0.863 0.861
BLEURT Scores 0.367 0.365 0.350 0.348 0.302 0.337 0.334
METEOR 0.198 0.194 0.153 0.142 0.153 0.159 0.173
ROUGE-2 0.041 0.037 0.034 0.033 0.029 0.034 0.037
@@
Atometric-F1 (M) 0.144 0.144 0.145 0.129 0.094 0.110 0.133
Atometric-F1 (S) 0.055 0.057 0.055 0.054 0.032 0.039 0.055
Atometric-P (M) 0.181 0.179 0.160 0.165 0.108 0.135 0.113
Atometric-R (M) 0.119 0.120 0.133 0.106 0.084 0.093 0.161
BERTScore F1 0.868 0.868 0.867 0.868 0.859 0.863 0.867
BLEURT Scores 0.340 0.345 0.341 0.339 0.286 0.323 0.340
METEOR 0.173 0.166 0.164 0.159 0.137 0.158 0.190
ROUGE-2 0.046 0.046 0.041 0.044 0.024 0.033 0.034
Table 4. Evaluation results of the candidate methods on the Map2Text task. The highest score(s) in each row are highlighted in bold and the second highest score(s) are in italic. Results of additional evaluation metrics (e.g., Atometric Precisions and Recalls) are provided in Appendix A.1. (L), (M), and (S) denote “loose”, “moderate”, and “strict” for strictness level. Note that the embedding inversion method relies on vec2text (Morris et al., 2023), which only supports the text-embedding-ada-002 model. As a result, no scores are provided for the Persona dataset.
@@
For the Persona and Red-Teaming Strategies datasets, CoT-RAG(1) consistently outperforms other AI methods across most metrics. This suggests that incorporating more complex reasoning structures, such as chain-of-thoughts, enhances the ability of LLMs to better align with the query position. Additionally, fine-tuning (FT) on these datasets shows clear improvements on Llama 3.1 compared to applying RAG(1) on its untuned version. This indicates that fine-tuning enables the LLM to better capture the specific nuances of the relation between the map locations and text.
@@
In contrast, for CS Research Ideas and CS Research Contexts, no single method clearly stands out. The best performer under most metrics is one of the GPT-4o based methods, indicating the importance of using a more capable foundation model. However, more complex prompting only leads to marginal improvements over the basic RAG(1) method. Interestingly, fine-tuning Llama 3.1 on these datasets appears to reduce its performance compared to the untuned RAG(1) model. One possible explanation is that these datasets contain highly specialized technical content, which may not align well with the general knowledge encoded in pre-trained LLMs. Without sufficient domain-specific data for fine-tuning, a model may lose some of its generalization capacity.
Additionally, the Embedding Inversion method demonstrates competitive performance, particularly under Atometric Recall, albeit with some loss in precision. On the Red-Teaming Strategies and CS Research Contexts, it achieves scores comparable to those of GPT-4o’s RAG(1) model, despite only employing a simple k-NN interpolation and a small pre-trained vec2text model (220M parameters) without additional training. This result highlights the potential of embedding inversion strategies for Map2Text tasks, when the originally embedding methods are accessible.
@@
Surprisingly, no candidate models surpass the EchoNearest baseline on the Red-Team Strategies dataset. This is likely due to the concentrated distribution of the dataset, as shown in Appendix Figure 7—with examples being so close to one another that distinguishing them becomes difficult. Intuitively, when the map is sparser (i.e., KNNs are farther apart), targeting relevance is more challenging. Conversely, when the map is locally denser (i.e., KNNs are closer together), relevance is easier to achieve, but innovation becomes harder, an issue that could be addressed in future work.
@@
Finally, the human baseline demonstrates superior performance compared to all candidate methods on the Persona dataset. Its primary advantage is the flexibility to navigate the visualization and extract high-level information, such as the distribution of topics across different areas and the directional relationships between nearest neighbors. This capability enables more accurate and contextually appropriate content generation, suggesting that there is still broad improvement space for better AI methods to incorporate similar high-level navigational strategies on the map.

### v2

@@ 3.3. Atometric: A New Evaluation Metric
Moderate The atomic statement can be logically inferred from the reference, or it can support the reference.
@@
Table 1. Overview of Atometric’s strictness levels. A bolded letter indicates each level’s abbreviation.
@@ 4. Experiments
With the evaluation setup, we are able to test several straightforward yet effective candidate methods for MapExplorer and compare them with a human baseline and a strong dummy baseline.
@@ 4.1. Candidate Methods
In this section, we detail the three potential solutions introduced in Section 2.2. Each solution represents a distinct approach to learning the mapping function $g$, which translates a low-dimensional query in ${\mathbb{R}}^{2}$ to the text space ${\mathcal{S}}$.
Direct Mapping.
The most straightforward approach is to directly learn the mapping $g:{\mathbb{R}}^{2}\to{\mathcal{S}}$ by fine-tuning (FT) a pre-trained model on each visualization map (therefore it is also computationally heavy). We fine-tune the 70B Llama 3.1 model using LoRA (Dubey et al., 2024), where the model takes a 2D coordinate as input and directly outputs the corresponding text. For instance, an input query would be, ”Convert the coordinate to text: [9.1054, 10.1339].”
Intermediate High-Dimensional Embedding.
Another method is to compose $g$ through two distinct mappings: $g_{1}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{H}$ and $g_{2}:{\mathbb{R}}^{H}\to{\mathcal{S}}$. For the first mapping, $g_{1}$, we use a k-NN interpolation algorithm to project the 2D coordinates into an $H$-dimensional text embedding space induced by a text embedding model. While more complex methods could be employed for $g_{1}$ (e.g., training a neural network), we select this straightforward approach as it requires no additional training and demonstrates comparable empirical performance in our preliminary experiments. The second mapping, $g_{2}$, then converts the high-dimensional embeddings into text using the vec2text model (Morris et al., 2023). Note that this approach heavily relies on the effectiveness of the pre-trained vec2text model, which may or may not be aligned with specific application domains.
Two-Stage Text Generation.
While LLMs excel at text generation, their zero-shot ability on visualization maps is inherently limited: maps encode spatial relationships that pre-trained LLMs cannot directly interpret. To bridge this gap, we design a retrieval-augmented framework that first grounds the LLM in the map’s context, then synthesizes text aligned with both the retrieved signals and the map’s latent structure. Specifically, for a query position, we retrieve texts from neighboring points in the visualization to construct a spatially informed context, which the LLM then uses to generate coherent, domain-aligned content. We formalize four variants of this approach, progressively refining how context is retrieved and utilized. For each query position ${\bm{x}}_{q}\in{\mathbb{R}}^{2}$,
@@
    1st-Order RAG: Retrieve texts associated with the query’s nearest neighbors $\mathcal{N}(\mathbf{x}_{q})$, then prompt the LLM to generate $\hat{s}$ using only this local context.
@@
    1st-Order RAG + Few-Shot: Extend (1) by adding pre-selected few-shot examples $\{(\mathcal{N}(\mathbf{x}_{i}),s_{i})\}_{i\in\mathcal{I}}$, where $\mathcal{I}$ is a fixed set of reference points. These examples teach the LLM to map neighborhoods into valid outputs.
@@
    2nd-Order RAG: Expand retrieval to second-order neighbors $\mathcal{N}(\mathcal{N}(\mathbf{x}_{q}))$ (2-hop connections). Unlike (2), few-shot examples $\{(\mathcal{N}(\mathbf{x}_{i}),s_{i})\}_{i\in\mathcal{I}}$ are not pre-selected but dynamically sourced from this extended neighborhood, ensuring regionally relevant context for generation.
@@
    1st-Order RAG + Chain-of-Thought: Augment (2) with step-by-step reasoning (e.g., ”First, identify themes in $\mathcal{N}(\mathbf{x}_{q})$…”) using chain-of-thought prompting (Wei et al., 2022). This explicates the LLM’s alignment of $\hat{s}$ with spatial context.
Implementation of each method can be found in Appendix E.
4.2. Datasets for Visualization
Dataset Name (Type) # Entries Avg. Len. Vis. Recipe
Persona (Text) 100k 27.10 gte-v1.5 + UMAP
Red Teaming Strategies (Text) 39k 84.06 ada-002 + UMAP
Research Idea (Text) 149k 41.57 ada-002 + LargeVis
Research Context (Text) 150k 40.96 ada-002 + LargeVis
Research Context (Network) 64k 45.35 LargeVis
Table 2. Basic statistics of the visualization maps. The average length is calculated using tiktoken.
We evaluate the candidate models on five diverse visualization maps to demonstrate the broad application of MapExplorer. The statistics of these datasets are summarized into Table 2. Note that the visualization pipelines for these datasets vary, which are hidden from candidate MapExplorer methods (except for embedding inversion), and a robust MapExplorer method should be capable of generalizing across different configurations.
Persona (Text)
We use an existing visualization [^2] https://huggingface.co/datasets/argilla/FinePersonas-v0.1-clustering-100k of a subset of the PersonHub dataset (Chan et al., 2024) where each text entry provides a brief description of a persona. MapExplorer can be applied to create diverse character profiles.
Red Teaming Strategies (Text)
We build a summarization dataset [^3] Details of the construction are provided in Appendix D.2. from the LLM Red Teaming dataset (Ganguli et al., 2022) which comprises human-generated and annotated red teaming dialogues aimed at identifying vulnerabilities in LLMs. Given a visualization map of these summaries, MapExplorer can be applied to generate unexplored red teaming strategies, enhancing the robustness testing of LLMs.
Research Ideas & Context (Text)
We generate two 2D maps from the MASSW dataset (Zhang et al., 2024), one representing research contexts (problems) and the other highlighting key ideas in computer science publications. An effective MapExplorer model can foster scientific innovation by producing new research problems or ideas for emerging topics from unexplored areas on the map.
Research Context (Network)
This dataset is similar to the Research Context (Text) dataset but includes only entries with sufficient citations. Instead of using text embeddings to map research contexts, we create a visualization based on the citation network. This network visualization represents a map creation process where high-dimensional text embedding is not an intermediate step.
@@ 4.3. Experiment Setup
Each dataset is divided into training and testing sets, as described in Section 3.1, with 200 samples consistently held out for testing across all datasets. We apply a random split for the Persona and Red Teaming Strategies datasets and a time-based split for the research publication datasets, where the most recent 200 entries (all in Dec 2023) are held out for testing, simulating the scenario of real-world research exploration.
@@
We prioritize our proposed evaluation metric, Atometric , as it is specifically designed to capture the nuanced semantic and contextual alignment. We report Atometric-F1 across three levels of strictness: loose (L), moderate (M), and strict (S), as well as Atometric-Precision (M) and Atometric-Recall (M) for a more comprehensive picture [^4] We use gpt-4o-2024-05-13 as the backbone LLM in Atometric.. In addition to Atometric, we include traditional evaluation metrics—BERTScore (F1) (Zhang et al., 2019), BLEURT (Sellam et al., 2020), METEOR (Lavie and Agarwal, 2007), and ROUGE-2 (Lin, 2004)—as reference points for comparison. Evaluation on additional metrics such as BLEU and other Atometric scores are provided in Appendix B.1.
@@ 4.4. Baselines
Metric Human EchoNearest
Persona (Text) Persona (Text) Red Teaming (Text) Idea (Text) Context (Text) Context (Network)
Atometric-F1 (L) 0.901 0.884 0.783 0.430 0.538 0.444
Atometric-F1 (M) 0.605 0.556 0.586 0.191 0.105 0.078
Atometric-F1 (S) 0.302 0.268 0.282 0.098 0.045 0.025
Atometric-P (M) 0.610 0.548 0.598 0.199 0.103 0.095
Atometric-R (M) 0.601 0.565 0.575 0.184 0.107 0.066
[…]
BLEURT Scores 0.431 0.454 0.436 0.325 0.323 0.293
METEOR 0.232 0.306 0.428 0.160 0.170 0.138
ROUGE-2 0.110 0.117 0.231 0.032 0.034 0.020
@@
We recruit human annotators to conduct the MapExplorer task using the Persona dataset, as this dataset does not require special domain expertise. Detailed settings for this study are provided in Appendix B.2. In brief, human annotators are given access to a partial visualization containing the query location and they are allowed to view the textual content of all visible data points. No constraints are imposed regarding their generation strategy, allowing them complete flexibility in navigating the visualization and generating content based on the map. We expect the human baseline to represent a strong benchmark compared to AI methods.
@@
For each query ${\bm{x}}_{q}$, the EchoNearest model simply retrieves and outputs the content of its closest neighbor in the 2D space. This naive baseline operates under the assumption that semantically similar content is clustered closely in the map, making the nearest neighbor a relevant choice for responding to the query, though it lacks novelty. As our gold-standard references all exhibit some level of novelty compared to their neighbors, we expect it to perform below both the human baseline and more advanced AI methods when evaluated using proper metrics.
@@
The evaluation results for the human and EchoNearest baselines are provided in Table 4.4. Several observations can be made.
First, under all Atometrics, the human baseline outperforms the EchoNearest model, as expected. However, under conventional metrics, the human baseline scores lower than the EchoNearest model. This discrepancy likely arises because conventional metrics emphasize lexical-level text similarity, favoring the EchoNearest model’s repetition of existing entries. In contrast, human responses, while more contextually relevant, tend to vary in language style and phrasing, which conventional metrics fail to capture. This finding validates the strength of Atometric in recognizing the semantic and contextual richness of human-generated content.
Second, the five datasets exhibit varying levels of complexity. The EchoNearest model performs relatively well on the Persona and Red Teaming Strategies datasets, suggesting that the text entries clustered closely in the visualization are highly similar, making these tasks easier to address. In contrast, the model’s performance on the other datasets is notably lower. This indicates that these tasks require generating more complex or novel content, which demands a deeper understanding of the research domain beyond simple proximity in the 2D space. Another interesting difference lies between two distinct visualization recipes for the same Research Context dataset: the performance of EchoNearest is better under the “text” setting, suggesting that visualization by text embedding may result in higher preservation of locality than by citation network.
In addition, the Atometric results across different levels of strictness reveal subtle distinctions between research contexts and research ideas, showcasing the strength of Atometric in capturing varying levels of sensitivity. At lower strictness levels, Atometric F1 (L) scores are higher for the Research Contexts. While each paper addresses a unique problem, the overarching topics are often related, leading to greater alignment in topic-level relevance. In contrast, at medium and high strictness levels (M and S), research ideas exhibit higher Atometric scores. This occurs because research ideas, while varied, frequently share concrete methodologies or designs, which stricter criteria capture more effectively. By distinguishing these nuanced differences, Atometric demonstrates its ability to assess content sensitivity at multiple levels.
@@ 4.5. Benchmark Results
Persona (Text) Atometric-F1 (L) 0.893 (0.012) 0.852 0.824 0.822 0.822 0.783 0.923
Atometric-F1 (M) 0.602 (0.019) 0.494 0.464 0.454 0.463 0.390 0.665
Atometric-F1 (S) 0.254 (0.013) 0.183 0.168 0.158 0.194 0.117 0.357
Atometric-P (M) 0.582 (0.019) 0.460 0.449 0.430 0.459 0.365 0.661
Atometric-R (M) 0.623 (0.019) 0.533 0.480 0.481 0.467 0.420 0.671
BERTScore F1 0.898 (0.001) 0.895 0.893 0.892 0.893 0.887 0.901
BLEURT Scores 0.474 (0.006) 0.454 0.441 0.433 0.406 0.416 0.495
METEOR 0.329 (0.009) 0.290 0.263 0.260 0.251 0.268 0.342
ROUGE-2 0.116 (0.006) 0.096 0.078 0.086 0.098 0.074 0.128
@@
Atometric-F1 (M) 0.514 (0.017) 0.480 0.451 0.414 0.497 0.437 0.428
Atometric-F1 (S) 0.202 (0.011) 0.200 0.168 0.144 0.185 0.167 0.126
Atometric-P (M) 0.479 (0.017) 0.445 0.397 0.358 0.508 0.384 0.329
Atometric-R (M) 0.555 (0.017) 0.521 0.520 0.491 0.486 0.507 0.612
BERTScore F1 0.904 (0.001) 0.902 0.901 0.897 0.894 0.898 0.877
BLEURT Scores 0.446 (0.004) 0.439 0.435 0.417 0.407 0.430 0.394
METEOR 0.433 (0.006) 0.424 0.412 0.371 0.390 0.405 0.312
ROUGE-2 0.250 (0.003) 0.250 0.244 0.232 0.203 0.235 0.114
@@
Atometric-F1 (M) 0.213 (0.010) 0.220 0.214 0.202 0.163 0.210 0.180
Atometric-F1 (S) 0.091 (0.005) 0.098 0.102 0.099 0.084 0.100 0.096
Atometric-P (M) 0.222 (0.010) 0.240 0.272 0.250 0.203 0.263 0.193
Atometric-R (M) 0.205 (0.010) 0.202 0.176 0.169 0.136 0.175 0.169
BERTScore F1 0.860 (0.001) 0.860 0.864 0.864 0.858 0.863 0.861
BLEURT Scores 0.367 (0.003) 0.365 0.350 0.348 0.302 0.337 0.334
METEOR 0.198 (0.004) 0.194 0.153 0.142 0.153 0.159 0.173
ROUGE-2 0.041 (0.002) 0.037 0.034 0.033 0.029 0.034 0.037
@@
Atometric-F1 (M) 0.144 (0.010) 0.144 0.145 0.129 0.094 0.110 0.133
Atometric-F1 (S) 0.055 (0.005) 0.057 0.055 0.054 0.032 0.039 0.055
Atometric-P (M) 0.181 (0.010) 0.179 0.160 0.165 0.108 0.135 0.113
Atometric-R (M) 0.119 (0.010) 0.120 0.133 0.106 0.084 0.093 0.161
BERTScore F1 0.868 (0.001) 0.868 0.867 0.868 0.859 0.863 0.867
BLEURT Scores 0.340 (0.004) 0.345 0.341 0.339 0.286 0.323 0.340
METEOR 0.173 (0.005) 0.166 0.164 0.159 0.137 0.158 0.190
ROUGE-2 0.046 (0.003) 0.046 0.041 0.044 0.024 0.033 0.034
[…]
Atometric-F1 (M) 0.083 (0.009) 0.058 0.055 0.046 0.037 0.071 0.068
Atometric-F1 (S) 0.025 (0.003) 0.022 0.017 0.008 0.011 0.021 0.021
Atometric-P (M) 0.102 (0.009) 0.072 0.067 0.072 0.048 0.065 0.062
Atometric-R (M) 0.070 (0.009) 0.048 0.048 0.034 0.031 0.077 0.076
BERTScore F1 0.862 (0.001) 0.860 0.861 0.860 0.854 0.813 0.858
BLEURT Scores 0.306 (0.004) 0.304 0.307 0.298 0.268 0.282 0.301
METEOR 0.131 (0.005) 0.115 0.118 0.104 0.125 0.169 0.149
ROUGE-2 0.025 (0.003) 0.022 0.018 0.021 0.012 0.015 0.018
Table 4. Evaluation results of the candidate methods on the MapExplorer task. The highest score(s) in each row are highlighted in bold and the second highest score(s) are underlined. Due to the space limit, only the standard error of the method CoT-RAG(1) is shown in parenthesis for reference.
@@
For the Persona and Red Teaming Strategies datasets, CoT-RAG(1) consistently outperforms other prompting methods. This suggests that incorporating more complex reasoning structures, such as chain of thoughts, enhances the ability of LLMs to better align with the query position. Additionally, fine-tuning (FT) on these datasets shows clear improvements on Llama 3.1 compared to applying RAG(1) on its untuned version. This indicates that fine-tuning enables the LLM to better capture the specific nuances of the relation between the map locations and text.
@@
In contrast, for the other three datasets on research papers, no single method clearly stands out. The best performer under most metrics is one of the GPT-4o based methods, indicating the importance of using a more capable foundation model. However, more complex prompting only leads to marginal improvements over the basic RAG(1) method. Interestingly, fine-tuning Llama 3.1 on these datasets appears to reduce its performance compared to the untuned RAG(1) model. One possible explanation is that these datasets contain highly specialized technical content, which may not align well with the general knowledge encoded in pre-trained LLMs. Without sufficient domain-specific data for fine-tuning, a model may lose some of its generalization capacity.
The Embedding Inversion method demonstrates competitive performance, particularly on the Persona dataset, which has the shortest average sequence length. It outperforms all pre-trained LLM-based methods despite relying solely on a simple k-NN interpolation and a small pre-trained vec2text model (220M parameters) without additional training. This result highlights the potential of embedding inversion strategies for MapExplorer tasks. However, as the generation length increases, the method struggles to produce coherent content, a limitation also observed in (Morris et al., 2023). We hypothesize that this is due to the inherent limitations of embedding inversion in handling logical structures.
@@
Surprisingly, no candidate models surpass the EchoNearest baseline on the Red Teaming Strategies dataset. This is likely due to the concentrated distribution of the dataset, as shown in Appendix Figure 9—with examples being so close to one another that distinguishing them becomes difficult. Intuitively, when the map is sparser (i.e., KNNs are farther apart), targeting relevance is more challenging. Conversely, when the map is locally denser (i.e., KNNs are closer together), relevance is easier to achieve, but innovation becomes harder, an issue that could be addressed in future work.
@@
Finally, the human baseline demonstrates superior performance compared to all candidate methods on the Persona dataset. Its primary advantage is the flexibility to navigate the visualization and extract high-level information, such as the distribution of topics across different areas and the directional relationships between nearest neighbors. This capability enables more accurate and contextually appropriate content generation, suggesting that there is still a broad improvement space for better AI methods to incorporate similar high-level navigational strategies on the map.
4.6. Interactive Demo
We offer an anonymous demo at https://mapexplorer-app.streamlit.app to complement our quantitative evaluation, enabling users to experiment with MapExplorer and test content generation at arbitrary locations. See Appendix A for screenshots and details [^5] Access to the demo may be subject to network conditions and local internet regulations. The demo is illustrative and not optimized for efficiency or usability. .